Circular No. 7350

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000D IN UGC 1767 G. M. Hurst, Basingstoke, England, reports the discovery by Tom

Boles, Wellingborough, of an apparent supernova (mag 16.2) on unfiltered CCD images taken on Jan. 22.928 and 24.806 UT in the course of the U.K. Nova/Supernova Patrol. SN 2000D is located at α = 2^{h}18^{m}04^{s}.84, δ = +38^{o}04′27′′.4 (equinox 2000.0; mean of two measures), which is 6′′ west and 11′′ south of the center of UGC 1767. Confirming images of SN 2000D were obtained by R. Arbour, South Wonston, on Jan. 24 and by M. Armstrong, Rolvenden, on Jan. 25.83. An image taken on 1999 Sept. 18.212 by Boles does not show the new object to a limit of mag ∼ 18. SN 2000D is not present on the second Palomar Sky Survey (limiting mag ∼ 19), where Hurst notes that the new star is slightly northwest of an H II region.

1E 161348–5055 G. P. Garmire, G. G. Pavlov, and A. B. Garmire, Pennsylvania State

University; and V. E. Zavlin, Max-Planck-Insitut für Extraterrestrische Physik, write: “The Chandra X-ray Observatory observed the supernova remnant (SNR) RCW 103 (G332.4–0.4) for 5.55 hr on 1999 Sept. 26. We have determined the following position of the pointlike source 1E 161348– 5055 seen near the center of RCW 103: α = 16^{h}17^{m}36^{s}.26, δ = −51^{o}02′25′′.0 (equinox 2000.0; uncertainty ∼ 0′′.2). No obvious optical counterpart (to mag R = 24) is seen closer than 3′′.95 from the source. The observed flux is 9.6 × 10^{−13} erg cm^{−2} s^{−1} in the band 0.4–10 keV, which corresponds to a luminosity of at least 1 × 10^{33} erg s^{−1} for a distance of 3300 pc. The light curve is compatible with a fragment of a sinusoid with period 5.97_{−0.44}^{+0.47} hr (90-percent confidence) and pulsed fraction ∼ 60 percent. We reanalyzed the ASCA archival data of this source obtained during a 16-hr period in Aug. 1993, and we found variations with period 6.05_{−0.35}^{+0.28} hr, the maximum of the first of the three observed peaks being at Aug. 18 ± 0.04 UT. The observed pulse fraction, 20–40 percent, is smaller than that found with Chandra, likely due to a larger contribution from the SNR background. The light curve of a 36-hr ASCA observation in Sept. 1997, when the source’s flux was an order of a magnitude lower, is consistent with a sinusoid with period 6.3 ± 0.6 hr. The 6-hr periodicity hints that the source may be a lowmass binary with an underluminous secondary component and unusually low x-ray luminosity. This is the first observation of such a long periodicity of a compact supernova remnant.”

2000 January 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green